OpenStudy (anonymous):

Medal Simplify 2x-3 [ x - (4-x) ]

OpenStudy (sdfgsdfgs):

start w the innermost bracket...what is x-(4-x) equal to?

OpenStudy (anonymous):

-4x - x

OpenStudy (misty1212):

HI!!

OpenStudy (sdfgsdfgs):

err...u are missing something there: what happens to the first x in the expression?

OpenStudy (misty1212):

first start with \[x-(4-x)\] distribute the minus sign get \[x-4+x\] combine like terms to get \[2x-4\] inside \[2x-3 [ x - (4-x) ]=2x-3(2x-4)\]

OpenStudy (sdfgsdfgs):

wait...no, please think about it again x-(4-x) there is no multiplication involved; just add and subtract....

OpenStudy (misty1212):

then with \[2x-3(2x-4)\] distribute the \(-3\) to get \[2x-6x+12\] then combine like terms

OpenStudy (anonymous):

yes how would i do this

OpenStudy (sdfgsdfgs):

@misty1212 i will let u take over :)

OpenStudy (misty1212):

\[2-6=-4\] so \[2x-6x=-4x\]

OpenStudy (anonymous):

okay how do i do this